Solve the following equations:
3s +12 = 0
Solve the following equations:
3s +12 = 0

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to subtract 12 to the both sides of the equation,
Then, we get,
= 3s + 12 - 12 = 0 - 12
= 3s = -12
Now,
We have to divide both sides of the equation by 3,
Then, we get,
= 3s/3 = -12/3
= s= - 4

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
3s + 12 = 0
3s = -12
S = -12/3
S = -4
 
Hope it helps!
Thank you ✌️